OpenStudy (anonymous):

help me pwetty pweese ! Solve each inequality. Write each set in interval notation. 7 < x + 6 *<=* 12

Nnesha (nnesha):

subtract 6 both side first

OpenStudy (anonymous):

Which would = 7<x*<=* 6 ?

Nnesha (nnesha):

both side left side 7-6 = ??

OpenStudy (anonymous):

1

OpenStudy (anonymous):

both sides of what ?

Nnesha (nnesha):

 when you have \[\le , \ge \] this sign use ] <--- this bracket and if you hve < or > this sign use parenthesis ) like this